### 16 Mar 2026 v1

# Open Biomedical Knowledge Graphs at Scale: Construction, Federation, and AI Agent Access with Samyama Graph Database

Abstract
Biomedical knowledge is fragmented across siloed databases—Reactome for pathways, STRING for protein interactions, Gene Ontology for functional annotations, ClinicalTrials.gov for study registries, and dozens more. Researchers routinely download flat files from each source and write bespoke scripts to cross-reference them, a process that is slow, error-prone, and not reproducible. We present two open-source biomedical knowledge graphs—Pathways KG (118,686 nodes, 834,785 edges from 5 sources) and Clinical Trials KG (7,774,446 nodes, 26,973,997 edges from 5 sources)—built on Samyama, a high-performance graph database written in Rust.
Our contributions are threefold. First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch Cypher loading, and portable snapshot export. Second, we demonstrate cross-KG federation: loading both snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “Which biological pathways are disrupted by drugs currently in Phase 3 trials for breast cancer?”—a query that neither KG can answer alone. Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries without manual tool authoring.
All data sources are open-license (CC BY 4.0, CC0, OBO). Snapshots, ETL code, and MCP configurations are publicly available. The combined federated graph (7.89M nodes, 27.8M edges) loads in 76 seconds on commodity hardware (Mac Mini M4, 16GB RAM), and the signature cross-KG query—“which pathways are disrupted by drugs in Phase 3 breast cancer trials?”—returns validated results in 2.1 seconds.
Keywords: Knowledge Graphs, Biomedical Data Integration, Graph Databases, Cross-KG Federation, Model Context Protocol, Clinical Trials, Biological Pathways, OpenCypher.
1 Introduction
Biological and clinical knowledge is distributed across dozens of public databases, each with its own schema, identifiers, and access patterns. A researcher investigating how a cancer drug affects cellular signaling must consult ClinicalTrials.gov for trial metadata, DrugBank for drug–target interactions, Reactome for pathway membership, STRING for protein–protein interactions, and Gene Ontology for functional annotations. These databases are maintained by independent communities, updated on different schedules, and stored in incompatible formats (TSV, XML, JSON, OBO, GAF).
Knowledge graphs (KGs) offer a natural integration point (Hogan et al., 2021): entities from different sources become nodes, and relationships between them become edges. A single Cypher query can traverse from a clinical trial to the biological pathways disrupted by its drug candidate—a query that would require joining across five separate databases in the flat-file paradigm.
However, constructing biomedical KGs at scale remains challenging. Prior efforts such as Bio2RDF (Callahan et al., 2013), Hetionet (Himmelstein et al., 2017), and the Clinical Knowledge Graph (Santos et al., 2022) have made significant progress, but each faces limitations: Bio2RDF requires a dedicated SPARQL endpoint infrastructure, Hetionet is a static dataset without an update pipeline, and the Clinical Knowledge Graph targets a single data source.
We address these limitations with three contributions:
  1. 1.
    Reproducible KG construction. We present ETL pipelines for two large-scale biomedical KGs—Pathways KG (5 sources, 119K nodes) and Clinical Trials KG (5 sources, 7.7M nodes)—built on Samyama Graph Database using a common pattern: download, parse, deduplicate, batch-load via Cypher, and export as portable .sgsnap snapshots.
  2. 2.
    Cross-KG federation. We show that loading multiple snapshots into a single graph tenant enables property-based federation—joining on shared identifiers (UniProt accessions, DrugBank IDs, NCBI Gene IDs) without ETL-time merging. We demonstrate five cross-KG query patterns spanning translational medicine.
  3. 3.
    Schema-driven AI agent access. Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration that auto-generates typed tools from the graph schema. LLM agents can query the KG via natural language without hand-coded tool definitions.
All code, snapshots, and MCP configurations are open-source [^1] Pathways KG: https://github.com/samyama-ai/pathways-kg; Clinical Trials KG: https://github.com/samyama-ai/clinicaltrials-kg; Snapshots: https://github.com/samyama-ai/samyama-graph/releases.
2 Background and Related Work
2.1 Biomedical Knowledge Graphs
Hetionet (Himmelstein et al., 2017) integrates 29 public resources into a single heterogeneous network (47K nodes, 2.25M edges, 24 edge types) for drug repurposing. While influential, it is a static dataset from 2017 with no update pipeline.
Bio2RDF (Callahan et al., 2013) converts life science databases to RDF and serves them via SPARQL endpoints. The RDF approach offers semantic interoperability but introduces complexity (OWL reasoning, blank nodes) and performance overhead for traversal-heavy queries.
Clinical Knowledge Graph (CKG) (Santos et al., 2022) builds a comprehensive biomedical graph from 25+ databases using Neo4j. CKG is the closest prior work to ours; we differ in three ways: (1) we use a Rust-native engine rather than Neo4j, achieving higher ingestion throughput; (2) we introduce portable snapshots for instant deployment; and (3) we provide schema-driven MCP servers for LLM agent integration.
PrimeKG (Chandak et al., 2023) constructs a precision medicine KG (129K nodes, 8M edges) from 20 sources. PrimeKG targets drug–disease prediction via graph neural networks; our focus is on interactive querying and AI agent access.
2.2 Graph Database Systems
Neo4j (Neo4j, Inc., 2024) is the dominant property graph database. Amazon Neptune, TigerGraph, and MemGraph offer alternatives with varying trade-offs. Samyama (Mandarapu and Kunkunuru, 2026) is a Rust-native graph database combining property graph storage, vector search, 22 metaheuristic solvers, and OpenCypher support in a single binary.
2.3 Model Context Protocol (MCP)
MCP (Anthropic, 2024) is an open standard for connecting LLM agents to external data sources via typed tools. An MCP server exposes a set of tools (functions with typed parameters and return values) that an LLM can invoke during a conversation. Schema-driven MCP generation—automatically creating tools from a KG schema—eliminates the manual tool authoring bottleneck.
3 Data Sources
Both KGs draw exclusively from open-license public databases. Table 1 summarizes the sources.
Table 1: Data sources for the two biomedical KGs. All are open-license and human-specific (organism 9606 where applicable).
KG Source Content License Size
Pathways Reactome Pathways, reactions, complexes CC BY 4.0 172 MB
STRING v12.0 Protein–protein interactions CC BY 4.0 900 MB
Gene Ontology GO terms & annotations OBO 265 MB
WikiPathways Community-curated pathways CC0 336 KB
UniProt Protein metadata, gene/disease/drug links CC BY 4.0 20 MB
Clinical ClinicalTrials.gov Trial registry (575K studies) Public domain API
MeSH (NLM) Disease hierarchy Free API
RxNorm (NLM) Drug normalization & ATC codes Free API
OpenFDA Adverse event reports (FAERS) Public domain API
PubMed (NLM) Trial-linked publications Free API
4 Knowledge Graph Construction
4.1 ETL Pattern
Both KGs follow a common five-phase pattern:
  1. 1.
    Download. A data downloader with resume support fetches all source files. Compressed files are decompressed automatically.
  2. 2.
    Parse & Filter. Source-specific parsers extract entities and relationships. Human-only filters are applied where applicable (organism ID 9606).
  3. 3.
    Deduplicate. A shared Registry tracks seen entities across phases (e.g., a protein loaded from Reactome is not re-created when encountered in STRING).
  4. 4.
    Batch Load. Nodes and edges are loaded via batched Cypher CREATE statements (50–100 entities per batch) through Samyama’s HTTP API.
  5. 5.
    Snapshot Export. The loaded graph is exported as a portable .sgsnap file (gzip JSON-lines), enabling instant restoration on any Samyama instance.
4.2 Pathways KG
The Pathways KG integrates molecular biology data into a graph with 5 node labels and 9 edge types (Table 2).
Table 2: Pathways KG schema.
Node Label Key Property Count
GOTerm go_id 51,897
Protein uniprot_id 37,990
Complex reactome_id 15,963
Reaction reactome_id 9,988
Pathway reactome_id 2,848
Total nodes 118,686
Edge Type Count
ANNOTATED_WITH 265,492
INTERACTS_WITH 227,818
PARTICIPATES_IN 140,153
CATALYZES 121,365
IS_A 58,799
COMPONENT_OF 8,186
PART_OF 7,122
REGULATES 2,986
CHILD_OF 2,864
Total edges 834,785
[…]
4.3 Clinical Trials KG
The Clinical Trials KG models the translational medicine domain with 15 node labels and 25 edge types (Table 3).
Table 3: Clinical Trials KG schema (abbreviated; 15 labels, 25 edge types).
Node Label Key Property Est. Count
ClinicalTrial nct_id 575,000+
Condition name, mesh_id varies
Intervention name, type varies
Drug rxnorm_cui, drugbank_id varies
Protein uniprot_id varies
Gene gene_id, symbol varies
MeSHDescriptor descriptor_id varies
Sponsor name, class varies
Site facility, country varies
Publication pmid, doi varies
AdverseEvent term, is_serious varies
ArmGroup label, type varies
Outcome measure, type varies
DrugClass atc_code, level varies
LabTest loinc_code varies
Total nodes 7,774,446
Total edges 26,973,997
The ETL pipeline queries the ClinicalTrials.gov API v2 for study metadata, enriches conditions with MeSH hierarchy, normalizes drug names via RxNorm, links adverse events from OpenFDA’s FAERS database, and associates publications from PubMed E-utilities.
5 Cross-KG Federation
5.1 Motivation
The Pathways KG knows molecular biology—which proteins interact, what pathways they participate in. The Clinical Trials KG knows translational medicine—which drugs are in trials, what conditions they treat. Neither alone can answer:
“Which biological pathways are disrupted by drugs currently in Phase 3 trials for breast cancer?”
[…]
5.2 Join Points
Three entity types appear in both KGs with matching identifiers:
Table 4: Cross-KG join points.
Entity Pathways Property Clinical Trials Property Identifier
Protein Protein.uniprot_id Protein.uniprot_id UniProt accession
Drug Drug.drugbank_id Drug.drugbank_id DrugBank ID
Gene Gene.gene_id Gene.gene_id NCBI Gene ID
5.3 Federation Mechanism
Samyama supports loading multiple snapshots into a single tenant. Each snapshot import appends nodes and edges to the existing graph. Since imports create new node IDs, entities from different snapshots with the same identifier (e.g., UniProt P04637 for TP53) exist as separate graph nodes. Cross-KG queries use property-based joins:
-- Pathways disrupted by Phase 3 breast cancer drugs
MATCH (ct:ClinicalTrial)-[:STUDIES]->(cond:Condition)
WHERE ct.phase = ’Phase3’ AND cond.name CONTAINS ’Breast’
WITH ct
MATCH (ct)-[:TESTS]->(int:Intervention)
-[:CODED_AS_DRUG]->(drug:Drug)
WITH DISTINCT drug
MATCH (drug)-[:TARGETS]->(prot1:Protein)
MATCH (prot2:Protein)-[:PARTICIPATES_IN]->(pw:Pathway)
WHERE prot1.uniprot_id = prot2.uniprot_id
RETURN pw.name AS pathway,
count(DISTINCT drug.name) AS drugs_targeting
ORDER BY drugs_targeting DESC LIMIT 15
The WHERE prot1.uniprot_id = prot2.uniprot_id clause is the cross-KG bridge—it joins a Protein node from the Clinical Trials snapshot with a Protein node from the Pathways snapshot.
5.4 Federation Query Patterns
We identify five cross-KG query patterns of increasing complexity:
Table 5: Cross-KG federation query patterns.
Pattern Clinical Trials $\rightarrow$ $\rightarrow$ Pathways
Drug $\rightarrow$ Pathway Trial $\rightarrow$ Drug (via intervention) Drug target $\rightarrow$ Protein $\rightarrow$ Pathway
Drug $\rightarrow$ GO Process Trial $\rightarrow$ Drug Drug target $\rightarrow$ Protein $\rightarrow$ GOTerm
Drug $\rightarrow$ PPI Network Trial $\rightarrow$ Drug $\rightarrow$ Protein target Target $\rightarrow$ INTERACTS_WITH neighbors
Disease $\rightarrow$ Pathway Gene $\rightarrow$ Disease association Gene $\rightarrow$ Protein $\rightarrow$ Pathway
Adverse Event $\rightarrow$ Pathway Trial $\rightarrow$ AdverseEvent $\rightarrow$ Drug Drug $\rightarrow$ Protein $\rightarrow$ Pathway
6 Schema-Driven MCP Server Generation
Each KG ships with a YAML configuration that defines domain-specific MCP tools. At startup, the MCP server:
  1. 1.
    Discovers schema from the running Samyama instance (GET /api/schema).
  2. 2.
    Auto-generates tools for each node label (search, get, count) and each edge type (find connections).
  3. 3.
    Registers domain tools from the YAML configuration—each tool is a parameterized Cypher template with typed inputs (e.g., protein_name: string, confidence_threshold: int).
The Pathways KG MCP server exposes 12 domain-specific tools (Table 6).
Table 6: Pathways KG MCP tools (subset).
Tool Description
pathway_members List proteins in a pathway (search by name)
interaction_partners PPI neighbors above confidence threshold
shared_pathways Pathways shared between two proteins
upstream_regulators Multi-hop PPI traversal up to N steps
drug_pathway_impact Pathways affected by a drug via protein targets
disease_pathways Pathways associated with a disease through gene links
go_enrichment GO terms enriched in pathway proteins
protein_function_summary Pathways, GO processes, disease associations for a protein
An LLM agent connected to this MCP server can answer “What pathways does TP53 participate in?” without writing Cypher. The agent calls pathway_members(protein_name="TP53") and receives structured results.
7 Evaluation
We evaluate construction time, query performance, and federation correctness on a Mac Mini M4 (Apple M4, 10 cores, 16GB RAM).
7.1 Construction Performance
Table 7: KG construction performance.
KG Nodes Edges Snapshot Size Import Time
Pathways KG 118,686 834,785 9 MB 1 s
Clinical Trials KG 7,774,446 26,973,997 711 MB 72 s
Drug enrichment (32 drugs) +33 +235 3 s
Combined 7,893,165 27,809,017 720 MB 76 s
Snapshot import uses gzip-compressed JSON-lines format. The Pathways KG loads in 1 second; the Clinical Trials KG, with its 7.8M nodes, loads in 72 seconds. A targeted drug enrichment step adds 32 breast cancer drug nodes with 36 TARGETS edges and 199 CODED_AS_DRUG edges in 3 seconds. The full pipeline—from empty tenant to federated graph—completes in 76 seconds.
7.2 Query Performance
We benchmark representative queries from each KG and the federated graph:
Table 8: Query latency on Mac Mini M4.
Query KG Results Latency
Label distribution (15 labels) Combined 15 rows 7.8 s
Proteins with uniprot_id Combined 1 row 0.5 s
Top pathways by protein count Pathways 10 rows 85 ms
PPI hub proteins (TP53: 739) Pathways 5 rows 1.4 s
Breast cancer trial count Clinical 5 rows 1.2 s
Drug $\rightarrow$ Pathway (6-hop) Combined 10 rows 2.1 s
The signature federated query—“pathways disrupted by breast cancer trial drugs”—traverses 6 hops across both KGs and returns in 2.1 seconds. Label distribution over the full 7.9M-node graph takes 7.8 seconds (full scan). Individual KG queries execute in under 1.5 seconds.
7.3 Federation Correctness
We validated cross-KG federation on the Mac Mini M4 with the following results:
  1. 1.
    Bridge construction: 32 breast cancer drug nodes were created from a curated DrugBank mapping, yielding 36 TARGETS edges (Drug $\rightarrow$ Protein) and 199 CODED_AS_DRUG edges (Intervention $\rightarrow$ Drug). The 1,544 matched drug interventions span 5,338 distinct drug names in breast cancer trials.
  2. 2.
    Join overlap: 37,990 Protein nodes with uniprot_id from the Pathways KG are available for joining. Of the 36 drug-target proteins, all are present in the Pathways KG’s protein-protein interaction network.
  3. 3.
    Federation query result: The 6-hop query “pathways disrupted by breast cancer trial drugs” returns 10 pathways, with Disease (21 drugs), Signal Transduction (19 drugs), Immune System (12 drugs), and Cell Cycle (12 drugs) as the top results. These align with known breast cancer biology: HER2 signaling (Trastuzumab), CDK4/6 cell cycle control (Palbociclib, Ribociclib), and immune checkpoint blockade (Pembrolizumab, Atezolizumab).
8 Interactive Visualization
Samyama Insight, a React-based frontend, provides schema-driven visualization of each KG:
  • Dashboard: Auto-generated panels showing label distribution, edge type counts, and property statistics per tenant.
  • Query Console: OpenCypher editor with result tables, JSON views, and EXPLAIN/PROFILE plan visualization.
  • Graph Simulation: Force-directed canvas with per-label colors/shapes, live activity particles, entity filtering, and a legend overlay. The simulation engine is fully schema-driven—it auto-configures from the tenant’s schema at runtime.
Demo recordings for both KGs are available as MP4 videos [^2] Cricket KG demo: https://github.com/samyama-ai/samyama-graph/releases/tag/kg-snapshots-v2; Pathways KG demo: https://github.com/samyama-ai/samyama-graph/releases/tag/kg-snapshots-v3.
9 Discussion
9.1 Property Joins vs. Entity Merging
Our federation approach uses property-based joins rather than merging entities at load time. This has trade-offs:
  • Advantages: Snapshots remain independent and composable; no load-order dependency; straightforward to add or remove a KG from the federation.
  • Disadvantages: Duplicate nodes inflate storage; property joins are slower than traversals on merged nodes; no referential integrity between the two copies of a Protein.
For production workloads, a post-load MERGE pass or ETL-time deduplication would eliminate duplicates. For exploration and prototyping—the primary use case for these KGs—property joins are pragmatic.
9.2 Limitations
  1. 1.
    Snapshot currency: Snapshots are point-in-time exports. Source databases (especially ClinicalTrials.gov) update continuously. Periodic re-export is required.
  2. 2.
    Identifier coverage: Not all proteins in the Clinical Trials KG have UniProt IDs (some have only gene symbols). The join overlap depends on identifier normalization quality.
  3. 3.
    Memory requirements: The combined 7.8M-node graph requires approximately 8GB RAM. Machines with less memory can load the Pathways KG alone (119K nodes, <<500MB).
  4. 4.
    No cross-tenant queries: Currently, federation requires loading both snapshots into a single tenant. Native cross-tenant query support is planned for a future Samyama release.
9.3 Generalizability
The pattern is not limited to biomedicine. Any domain with shared identifiers across data sources can use the same approach: build independent KGs with common entity properties, export as snapshots, load into a single tenant, and query across them. We have applied this pattern to sports (Cricket KG, 36K nodes) and industrial operations (AssetOps KG, 781 nodes) in addition to the biomedical KGs described here.
10 Conclusion
We presented two open-source biomedical knowledge graphs—Pathways KG and Clinical Trials KG—built on Samyama Graph Database. Together they contain 7.89 million nodes and 27.8 million edges from 10 public data sources. We demonstrated that loading both snapshots into a single graph tenant, combined with a targeted drug-target enrichment step (32 drugs, 36 TARGETS edges), enables cross-KG federated queries that bridge molecular biology and translational medicine. The 6-hop query “pathways disrupted by Phase 3 breast cancer drugs” returns biologically validated results—Signal Transduction, Immune System, Cell Cycle—in 2.1 seconds on commodity hardware. We introduced schema-driven MCP server generation for LLM agent access to graph queries.
All code, snapshots, ETL pipelines, and MCP configurations are open-source under Apache License 2.0. Researchers can reproduce the federation—from snapshot import to cross-KG queries—in under 2 minutes on commodity hardware.
Data and Code Availability

### 17 Mar 2026 v2

@@
 Abstract
-Biomedical knowledge is fragmented across siloed databases—Reactome for pathways, STRING for protein interactions, Gene Ontology for functional annotations, ClinicalTrials.gov for study registries, and dozens more.
+Biomedical knowledge is fragmented across siloed databases—Reactome for pathways, STRING for protein interactions, Gene Ontology for functional annotations, ClinicalTrials.gov for study registries, DrugBank for drug vocabularies, DGIdb for drug–gene interactions, SIDER for side effects, and dozens more.
 Researchers routinely download flat files from each source and write bespoke scripts to cross-reference them, a process that is slow, error-prone, and not reproducible.
-We present two open-source biomedical knowledge graphs—Pathways KG (118,686 nodes, 834,785 edges from 5 sources) and Clinical Trials KG (7,774,446 nodes, 26,973,997 edges from 5 sources)—built on Samyama, a high-performance graph database written in Rust.
+We present three open-source biomedical knowledge graphs—Pathways KG (118,686 nodes, 834,785 edges from 5 sources), Clinical Trials KG (7,774,446 nodes, 26,973,997 edges from 5 sources), and Drug Interactions KG (32,726 nodes, 191,970 edges from 3 sources)—built on Samyama, a high-performance graph database written in Rust.
 Our contributions are threefold.
-First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch Cypher loading, and portable snapshot export.
-Second, we demonstrate cross-KG federation: loading both snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “Which biological pathways are disrupted by drugs currently in Phase 3 trials for breast cancer?”—a query that neither KG can answer alone.
-Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries without manual tool authoring.
-All data sources are open-license (CC BY 4.0, CC0, OBO).
+First, we describe a reproducible ETL pattern for constructing large-scale KGs from heterogeneous public data sources, with cross-source deduplication, batch loading (both Python Cypher and Rust native loaders), and portable snapshot export.
+Second, we demonstrate cross-KG federation: loading all three snapshots into a single graph tenant enables property-based joins across datasets, answering questions like “For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”—a query that no single KG can answer alone.
+Third, we introduce schema-driven MCP server generation: each KG automatically exposes typed tools for LLM agents via the Model Context Protocol, enabling natural-language access to graph queries.
+We evaluate domain-specific MCP tools against text-to-Cypher and standalone GPT-4o on a new BiomedQA benchmark (40 pharmacology questions), achieving 98% accuracy vs.
+0% for text-to-Cypher and 75% for standalone GPT-4o.
+All data sources are open-license (CC BY 4.0, CC0, OBO, public domain).
 Snapshots, ETL code, and MCP configurations are publicly available.
-The combined federated graph (7.89M nodes, 27.8M edges) loads in 76 seconds on commodity hardware (Mac Mini M4, 16GB RAM), and the signature cross-KG query—“which pathways are disrupted by drugs in Phase 3 breast cancer trials?”—returns validated results in 2.1 seconds.
-Keywords: Knowledge Graphs, Biomedical Data Integration, Graph Databases, Cross-KG Federation, Model Context Protocol, Clinical Trials, Biological Pathways, OpenCypher.
+The combined federated graph (7.9M nodes, 28M edges) loads in approximately 3 minutes from portable snapshots on commodity cloud hardware (AWS g4dn.4xlarge, 62 GB RAM), and cross-KG queries complete in 80 ms–4 s.
+Keywords: Knowledge Graphs, Biomedical Data Integration, Graph Databases, Cross-KG Federation, Model Context Protocol, Clinical Trials, Biological Pathways, Drug Interactions, Pharmacogenomics, OpenCypher.
 1 Introduction
@@
 However, constructing biomedical KGs at scale remains challenging.
-Prior efforts such as Bio2RDF (Callahan et al., 2013), Hetionet (Himmelstein et al., 2017), and the Clinical Knowledge Graph (Santos et al., 2022) have made significant progress, but each faces limitations: Bio2RDF requires a dedicated SPARQL endpoint infrastructure, Hetionet is a static dataset without an update pipeline, and the Clinical Knowledge Graph targets a single data source.
+Prior efforts such as Bio2RDF (Belleau et al., 2008), Hetionet (Himmelstein et al., 2017), and the Clinical Knowledge Graph (Santos et al., 2022) have made significant progress, but each faces limitations: Bio2RDF requires a dedicated SPARQL endpoint infrastructure, Hetionet is a static dataset without an update pipeline, and the Clinical Knowledge Graph targets a single data source.
 We address these limitations with three contributions:
@@
     Reproducible KG construction.
-We present ETL pipelines for two large-scale biomedical KGs—Pathways KG (5 sources, 119K nodes) and Clinical Trials KG (5 sources, 7.7M nodes)—built on Samyama Graph Database using a common pattern: download, parse, deduplicate, batch-load via Cypher, and export as portable .sgsnap snapshots.
+We present ETL pipelines for three biomedical KGs—Pathways KG (5 sources, 119K nodes), Clinical Trials KG (5 sources, 7.7M nodes), and Drug Interactions KG (3 sources, 33K nodes)—built on Samyama Graph Database using a common pattern: download, parse, deduplicate, load (Python Cypher or Rust native), and export as portable .sgsnap snapshots.
+The Rust native loader creates the Drug Interactions KG (32K nodes, 192K edges) in under 1 second.
   2. 2.
     Cross-KG federation.
-We show that loading multiple snapshots into a single graph tenant enables property-based federation—joining on shared identifiers (UniProt accessions, DrugBank IDs, NCBI Gene IDs) without ETL-time merging.
-We demonstrate five cross-KG query patterns spanning translational medicine.
+We show that loading multiple snapshots into a single graph tenant enables property-based federation—joining on shared identifiers (UniProt accessions, DrugBank IDs, gene names) without ETL-time merging.
+We demonstrate cross-KG query patterns spanning molecular biology, translational medicine, and pharmacogenomics.
   3. 3.
     Schema-driven AI agent access.
-Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration that auto-generates typed tools from the graph schema.
-LLM agents can query the KG via natural language without hand-coded tool definitions.
-All code, snapshots, and MCP configurations are open-source [^1] Pathways KG: https://github.com/samyama-ai/pathways-kg; Clinical Trials KG: https://github.com/samyama-ai/clinicaltrials-kg; Snapshots: https://github.com/samyama-ai/samyama-graph/releases.
+Each KG ships with a Model Context Protocol (MCP) (Anthropic, 2024) server configuration with domain-specific tools.
+We evaluate MCP tools against text-to-Cypher (GPT-4o) and standalone GPT-4o on a new BiomedQA benchmark: MCP tools achieve 98% accuracy (39/40) vs.
+0% for text-to-Cypher and 75% for standalone GPT-4o.
+All code, snapshots, and MCP configurations are open-source [^1] Pathways KG: https://github.com/samyama-ai/pathways-kg; Clinical Trials KG: https://github.com/samyama-ai/clinicaltrials-kg; Drug Interactions KG: https://github.com/samyama-ai/druginteractions-kg; Snapshots: https://github.com/samyama-ai/samyama-graph/releases.
 2 Background and Related Work
@@
 While influential, it is a static dataset from 2017 with no update pipeline.
-Bio2RDF (Callahan et al., 2013) converts life science databases to RDF and serves them via SPARQL endpoints.
+Bio2RDF (Belleau et al., 2008) converts life science databases to RDF and serves them via SPARQL endpoints.
 The RDF approach offers semantic interoperability but introduces complexity (OWL reasoning, blank nodes) and performance overhead for traversal-heavy queries.
@@
 Both KGs draw exclusively from open-license public databases. Table 1 summarizes the sources.
-Table 1: Data sources for the two biomedical KGs. All are open-license and human-specific (organism 9606 where applicable).
+Table 1: Data sources for the three biomedical KGs. All are open-license and human-specific (organism 9606 where applicable).
 KG Source Content License Size
@@
 PubMed (NLM) Trial-linked publications Free API
+Drug Int. DrugBank CC0 Drug vocabulary (19K drugs) CC0 3 MB
+DGIdb Drug–gene interactions (38K edges) Open 24 MB
+SIDER Side effects & indications $C$-BY-SA 22 MB
 4 Knowledge Graph Construction
@@
 The ETL pipeline queries the ClinicalTrials.gov API v2 for study metadata, enriches conditions with MeSH hierarchy, normalizes drug names via RxNorm, links adverse events from OpenFDA’s FAERS database, and associates publications from PubMed E-utilities.
+4.4 Drug Interactions KG
+The Drug Interactions KG models drug–gene interactions, side effects, and indications with 4 node labels and 3 edge types (Table 4).
+Table 4: Drug Interactions KG schema.
+Node Label Key Property Count
+Drug drugbank_id 19,842
+Gene gene_name 4,182
+SideEffect meddra_id 5,858
+Indication meddra_id 2,844
+Total nodes 32,726
+Edge Type Count
+INTERACTS_WITH_GENE 38,033
+HAS_SIDE_EFFECT 139,193
+HAS_INDICATION 14,744
+Total edges 191,970
+The ETL consists of two phases.
+Phase 1 (DrugBank + DGIdb): 19,842 Drug nodes are created from DrugBank CC0 vocabulary CSV, with all synonyms indexed for cross-source name matching (52,154 synonyms).
+DGIdb interactions yield 4,182 Gene nodes and 38,033 INTERACTS_WITH_GENE edges with interaction types (inhibitor, activator, etc.).
+Phase 2 (SIDER): STITCH compound IDs are mapped to DrugBank names via synonym lookup, yielding 5,858 SideEffect nodes with 139,193 HAS_SIDE_EFFECT edges and 2,844 Indication nodes with 14,744 HAS_INDICATION edges.
+A Rust native loader (druginteractions_loader.rs) performs the full ETL in 928 ms using direct GraphStore API calls (no Cypher parsing), compared to $\sim$}50 minutes via the Python HTTP loader.
 5 Cross-KG Federation
@@
 The Clinical Trials KG knows translational medicine—which drugs are in trials, what conditions they treat.
-Neither alone can answer:
-“Which biological pathways are disrupted by drugs currently in Phase 3 trials for breast cancer?”
+The Drug Interactions KG knows pharmacology—which genes a drug targets, what side effects it causes.
+No single KG can answer:
+“For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”
 […]
 5.2 Join Points
-Three entity types appear in both KGs with matching identifiers:
-Table 4: Cross-KG join points.
-Entity Pathways Property Clinical Trials Property Identifier
-Protein Protein.uniprot_id Protein.uniprot_id UniProt accession
-Drug Drug.drugbank_id Drug.drugbank_id DrugBank ID
-Gene Gene.gene_id Gene.gene_id NCBI Gene ID
+Shared entity types with matching identifiers enable cross-KG joins:
+Table 5: Cross-KG join points across the three biomedical KGs.
+Entity From KG To KG Join Property Identifier
+Gene/Protein Drug Int. Pathways gene_name = name Gene symbol
+Drug Drug Int. Clin. Trials name = Intervention.name Drug name
+Drug Drug Int. Clin. Trials drugbank_id DrugBank ID
+Protein Clin. Trials Pathways uniprot_id UniProt accession
 5.3 Federation Mechanism
@@
 7 Evaluation
-We evaluate construction time, query performance, and federation correctness on a Mac Mini M4 (Apple M4, 10 cores, 16GB RAM).
+We evaluate construction time, query performance, and AI agent access on an AWS g4dn.4xlarge instance (16 vCPU AMD EPYC, 62 GB RAM, NVIDIA A10G GPU).
 7.1 Construction Performance
-Table 7: KG construction performance.
-KG Nodes Edges Snapshot Size Import Time
-Pathways KG 118,686 834,785 9 MB 1 s
-Clinical Trials KG 7,774,446 26,973,997 711 MB 72 s
-Drug enrichment (32 drugs) +33 +235 3 s
-Combined 7,893,165 27,809,017 720 MB 76 s
+Table 8: KG construction performance (snapshot import on AWS g4dn.4xlarge).
+KG Nodes Edges Snapshot Import Rust ETL
+Pathways KG 118,686 834,785 9 MB 3.4 s
+Drug Interactions KG 32,726 191,970 1.9 MB 0.7 s 0.9 s
+Clinical Trials KG 7,774,446 26,973,997 711 MB 177 s
+Combined 7,925,858 28,000,752 722 MB 181 s
 Snapshot import uses gzip-compressed JSON-lines format.
-The Pathways KG loads in 1 second; the Clinical Trials KG, with its 7.8M nodes, loads in 72 seconds.
-A targeted drug enrichment step adds 32 breast cancer drug nodes with 36 TARGETS edges and 199 CODED_AS_DRUG edges in 3 seconds.
-The full pipeline—from empty tenant to federated graph—completes in 76 seconds.
+The Pathways KG loads in 3.6 seconds; the Drug Interactions KG in 0.7 seconds; and the Clinical Trials KG, with its 7.8M nodes, loads in 174 seconds.
+The Drug Interactions KG also has a Rust native loader that constructs the full graph from source files (DrugBank CSV, DGIdb TSV, SIDER TSV) in 928 ms—orders of magnitude faster than the Python HTTP-based ETL.
+The combined federated graph (7.9M nodes, 28M edges) uses 33 GB RAM (53% of 62 GB available on the AWS instance).
 7.2 Query Performance
 We benchmark representative queries from each KG and the federated graph:
-Table 8: Query latency on Mac Mini M4.
+Table 9: Query latency on AWS g4dn.4xlarge (7.9M nodes loaded).
 Query KG Results Latency
-Label distribution (15 labels) Combined 15 rows 7.8 s
-Proteins with uniprot_id Combined 1 row 0.5 s
-Top pathways by protein count Pathways 10 rows 85 ms
-PPI hub proteins (TP53: 739) Pathways 5 rows 1.4 s
-Breast cancer trial count Clinical 5 rows 1.2 s
-Drug $\rightarrow$ Pathway (6-hop) Combined 10 rows 2.1 s
-The signature federated query—“pathways disrupted by breast cancer trial drugs”—traverses 6 hops across both KGs and returns in 2.1 seconds.
-Label distribution over the full 7.9M-node graph takes 7.8 seconds (full scan).
-Individual KG queries execute in under 1.5 seconds.
+Drug gene targets (Metformin) Drug Int. 5 rows 83 ms
+Shared gene targets (2 drugs) Drug Int. 1 row 84 ms
+Top pathways by protein count Pathways 10 rows 97 ms
+PPI partners (TP53: Q9UQ61) Pathways 10 rows 96 ms
+Side effects of Warfarin Drug Int. 10 rows 82 ms
+Drugs for diabetes $\rightarrow$ gene targets Drug+Path 10 rows 3.6 s
+EGFR drugs $\rightarrow$ side effects Drug Int. 10 rows 3.7 s
+Cross-KG: PTGS1 drugs + side effects All 3 10 rows 3.7 s
+Simple single-KG queries (drug lookups, gene targets, side effects) complete in 80–100 ms.
+Multi-hop cross-KG joins (e.g., “drugs targeting PTGS1 and their most common side effects”) complete in 3–4 s on the full 7.9M-node graph.
+7.3 BiomedQA Benchmark
+We introduce BiomedQA, a benchmark of 40 pharmacology questions across 7 categories over the three federated KGs.
+We compare three approaches: domain-specific MCP tools (parameterized Cypher templates), text-to-Cypher (GPT-4o generates Cypher from a schema prompt with 3 examples), and standalone GPT-4o (no database access).
+Table 10: BiomedQA results (40 questions, 7.9M nodes, 3 federated KGs).
+Approach Accuracy Avg Latency Avg Tokens
+GPT-4o standalone 30/40 (75%) 2,474 ms 213
+Text-to-Cypher (GPT-4o) 0/40 (0%) 986 ms 548
+MCP tools 39/40 (98%) 651 ms 0
+The text-to-Cypher result (0%) is striking: GPT-4o generates syntactically plausible Cypher that silently returns empty results due to schema mismatches (hallucinated property names, incorrect multi-MATCH patterns).
+MCP tools eliminate this class of failure entirely.
+The benchmark and evaluation scripts are open-source.
 7.4 Federation Correctness
@@
 10 Conclusion
-We presented two open-source biomedical knowledge graphs—Pathways KG and Clinical Trials KG—built on Samyama Graph Database.
-Together they contain 7.89 million nodes and 27.8 million edges from 10 public data sources.
-We demonstrated that loading both snapshots into a single graph tenant, combined with a targeted drug-target enrichment step (32 drugs, 36 TARGETS edges), enables cross-KG federated queries that bridge molecular biology and translational medicine.
-The 6-hop query “pathways disrupted by Phase 3 breast cancer drugs” returns biologically validated results—Signal Transduction, Immune System, Cell Cycle—in 2.1 seconds on commodity hardware.
-We introduced schema-driven MCP server generation for LLM agent access to graph queries.
-All code, snapshots, ETL pipelines, and MCP configurations are open-source under Apache License 2.0.
-Researchers can reproduce the federation—from snapshot import to cross-KG queries—in under 2 minutes on commodity hardware.
+We presented three open-source biomedical knowledge graphs—Pathways KG (119K nodes from 5 sources), Clinical Trials KG (7.7M nodes from 5 sources), and Drug Interactions KG (33K nodes from 3 sources)—built on Samyama Graph Database.
+Together they contain 7.9 million nodes and 28 million edges from 13 public data sources.
+Loading all three snapshots into a single graph tenant enables cross-KG federated queries that bridge molecular biology, translational medicine, and pharmacogenomics.
+Cross-KG queries complete in 80 ms–4 s on commodity hardware (62 GB RAM).
+A Rust native loader constructs the Drug Interactions KG in under 1 second.
+We introduced domain-specific MCP tools for LLM agent access and evaluated them on a new BiomedQA benchmark (40 pharmacology questions): MCP tools achieve 98% accuracy vs.
+0% for text-to-Cypher and 75% for standalone GPT-4o, demonstrating that for domain-specific data access, the LLM’s role should be tool selection and argument extraction, not query generation.
+All code, snapshots, ETL pipelines, benchmarks, and MCP configurations are open-source. Researchers can reproduce the full 3-KG federation—from snapshot import to cross-KG queries—in under 3 minutes.
 Data and Code Availability

### 18 Mar 2026 v3

@@
 We evaluate domain-specific MCP tools against text-to-Cypher and standalone GPT-4o on a new BiomedQA benchmark (40 pharmacology questions), achieving 98% accuracy vs.
-0% for text-to-Cypher and 75% for standalone GPT-4o.
+85% for schema-aware text-to-Cypher and 75% for standalone GPT-4o, with zero schema errors.
 All data sources are open-license (CC BY 4.0, CC0, OBO, public domain).
 Snapshots, ETL code, and MCP configurations are publicly available.
-The combined federated graph (7.9M nodes, 28M edges) loads in approximately 3 minutes from portable snapshots on commodity cloud hardware (AWS g4dn.4xlarge, 62 GB RAM), and cross-KG queries complete in 80 ms–4 s.
+The combined federated graph (7.9M nodes, 28M edges) loads in approximately 3 minutes from portable snapshots on commodity cloud hardware (62 GB RAM), with single-KG queries completing in 80–100 ms and cross-KG federation joins in 1–4 s.
 Keywords: Knowledge Graphs, Biomedical Data Integration, Graph Databases, Cross-KG Federation, Model Context Protocol, Clinical Trials, Biological Pathways, Drug Interactions, Pharmacogenomics, OpenCypher.
@@
 We evaluate MCP tools against text-to-Cypher (GPT-4o) and standalone GPT-4o on a new BiomedQA benchmark: MCP tools achieve 98% accuracy (39/40) vs.
-0% for text-to-Cypher and 75% for standalone GPT-4o.
+85% for schema-aware text-to-Cypher and 75% for standalone GPT-4o, with zero schema errors.
 All code, snapshots, and MCP configurations are open-source [^1] Pathways KG: https://github.com/samyama-ai/pathways-kg; Clinical Trials KG: https://github.com/samyama-ai/clinicaltrials-kg; Drug Interactions KG: https://github.com/samyama-ai/druginteractions-kg; Snapshots: https://github.com/samyama-ai/samyama-graph/releases.
@@
 3 Data Sources
-Both KGs draw exclusively from open-license public databases. Table 1 summarizes the sources.
+All three KGs draw exclusively from open-license public databases. Table 1 summarizes the sources.
 Table 1: Data sources for the three biomedical KGs. All are open-license and human-specific (organism 9606 where applicable).
@@
 4.1 ETL Pattern
-Both KGs follow a common five-phase pattern:
+All three KGs follow a common five-phase pattern:
   1. 1.
@@
 “For drugs indicated for diabetes, what are their gene targets and which biological pathways do those targets participate in?”
-[…]
+This query requires traversing three KGs: Drug $\xrightarrow{\text{HAS\_INDICATION}}$ Indication (Drug Interactions KG), Drug $\xrightarrow{\text{INTERACTS\_WITH\_GENE}}$ Gene $\xrightarrow{\text{bridge: gene\_name = name}}$ Protein $\xrightarrow{\text{PARTICIPATES\_IN}}$ Pathway (Pathways KG).
+The WHERE p.name = g.gene_name clause bridges Drug Interactions to Pathways.
 5.2 Join Points
@@
 Cross-KG queries use property-based joins:
--- Pathways disrupted by Phase 3 breast cancer drugs
-MATCH (ct:ClinicalTrial)-[:STUDIES]->(cond:Condition)
-WHERE ct.phase = ’Phase3’ AND cond.name CONTAINS ’Breast’
-WITH ct
-MATCH (ct)-[:TESTS]->(int:Intervention)
--[:CODED_AS_DRUG]->(drug:Drug)
-WITH DISTINCT drug
-MATCH (drug)-[:TARGETS]->(prot1:Protein)
-MATCH (prot2:Protein)-[:PARTICIPATES_IN]->(pw:Pathway)
-WHERE prot1.uniprot_id = prot2.uniprot_id
-RETURN pw.name AS pathway,
-count(DISTINCT drug.name) AS drugs_targeting
-ORDER BY drugs_targeting DESC LIMIT 15
-The WHERE prot1.uniprot_id = prot2.uniprot_id clause is the cross-KG bridge—it joins a Protein node from the Clinical Trials snapshot with a Protein node from the Pathways snapshot.
+-- Drug targets -> biological pathways (Drug Interactions -> Pathways)
+MATCH (d:Drug {name: ’Metformin’})-[:INTERACTS_WITH_GENE]->(g:Gene)
+MATCH (p:Protein)-[:PARTICIPATES_IN]->(pw:Pathway)
+WHERE p.name = g.gene_name
+RETURN g.gene_name, pw.name LIMIT 10
+-- Drug -> clinical trials testing it (Drug Interactions -> Clinical Trials)
+MATCH (d:Drug {name: ’Warfarin’})
+MATCH (i:Intervention)<-[:TESTS]-(ct:ClinicalTrial)
+WHERE i.name = d.name
+RETURN ct.nct_id, ct.phase LIMIT 10
+-- Breast cancer trial landscape (Clinical Trials only)
+MATCH (ct:ClinicalTrial)-[:STUDIES]->(c:Condition)
+WHERE c.name CONTAINS ’Breast’
+RETURN c.name, count(ct) AS trials
+ORDER BY trials DESC LIMIT 5
+The WHERE p.name = g.gene_name clause is the Drug Interactions $\to$ Pathways bridge—it joins a Gene node from the Drug Interactions KG with a Protein node from the Pathways KG using gene symbol as the shared identifier.
+The WHERE i.name = d.name clause bridges Drug Interactions $\to$ Clinical Trials via drug name.
 5.4 Federation Query Patterns
@@
 Table 6: Cross-KG federation query patterns.
-Pattern Clinical Trials $\rightarrow$ $\rightarrow$ Pathways
-Drug $\rightarrow$ Pathway Trial $\rightarrow$ Drug (via intervention) Drug target $\rightarrow$ Protein $\rightarrow$ Pathway
-Drug $\rightarrow$ GO Process Trial $\rightarrow$ Drug Drug target $\rightarrow$ Protein $\rightarrow$ GOTerm
-Drug $\rightarrow$ PPI Network Trial $\rightarrow$ Drug $\rightarrow$ Protein target Target $\rightarrow$ INTERACTS_WITH neighbors
-Disease $\rightarrow$ Pathway Gene $\rightarrow$ Disease association Gene $\rightarrow$ Protein $\rightarrow$ Pathway
-Adverse Event $\rightarrow$ Pathway Trial $\rightarrow$ AdverseEvent $\rightarrow$ Drug Drug $\rightarrow$ Protein $\rightarrow$ Pathway
+Pattern From KG Bridge To KG
+Drug $\rightarrow$ Pathway Drug Int. (Gene) gene_name Pathways (Protein)
+Drug $\rightarrow$ Trial Drug Int. (Drug) name Clin. Trials (Intervention)
+Drug $\rightarrow$ GO terms Drug Int. (Gene) gene_name Pathways (GOTerm)
+Drug SE $\rightarrow$ Pathway Drug Int. (SE+Gene) gene_name Pathways (Protein)
+Trial $\rightarrow$ Side effects Clin. Trials (Intervention) name Drug Int. (Drug)
 6 Schema-Driven MCP Server Generation
@@
     Registers domain tools from the YAML configuration—each tool is a parameterized Cypher template with typed inputs (e.g., protein_name: string, confidence_threshold: int).
-The Pathways KG MCP server exposes 12 domain-specific tools (Table 7).
+Each KG’s MCP server exposes domain-specific tools.
+The Pathways KG has 12 tools (Table 7), the Drug Interactions KG has 12 tools (e.g., drug_interactions, interaction_checker, polypharmacy_risk, drug_side_effects), and the Clinical Trials KG has 15 tools.
 Table 7: Pathways KG MCP tools (subset).
@@
 Snapshot import uses gzip-compressed JSON-lines format.
-The Pathways KG loads in 3.6 seconds; the Drug Interactions KG in 0.7 seconds; and the Clinical Trials KG, with its 7.8M nodes, loads in 174 seconds.
+The Pathways KG loads in 3.4 seconds; the Drug Interactions KG in 0.7 seconds; and the Clinical Trials KG, with its 7.8M nodes, loads in 177 seconds.
 The Drug Interactions KG also has a Rust native loader that constructs the full graph from source files (DrugBank CSV, DGIdb TSV, SIDER TSV) in 928 ms—orders of magnitude faster than the Python HTTP-based ETL.
@@
 Side effects of Warfarin Drug Int. 10 rows 82 ms
-Drugs for diabetes $\rightarrow$ gene targets Drug+Path 10 rows 3.6 s
-EGFR drugs $\rightarrow$ side effects Drug Int. 10 rows 3.7 s
-Cross-KG: PTGS1 drugs + side effects All 3 10 rows 3.7 s
+Diabetes drugs $\rightarrow$ pathways Drug Int.+Path 5 rows 3.9 s
+Warfarin $\rightarrow$ trial conditions Drug Int.+CT 5 rows 5.8 s
+PTGS1 drugs $\rightarrow$ side effects Drug Int. 10 rows 3.7 s
 Simple single-KG queries (drug lookups, gene targets, side effects) complete in 80–100 ms.
@@
 We introduce BiomedQA, a benchmark of 40 pharmacology questions across 7 categories over the three federated KGs.
-We compare three approaches: domain-specific MCP tools (parameterized Cypher templates), text-to-Cypher (GPT-4o generates Cypher from a schema prompt with 3 examples), and standalone GPT-4o (no database access).
+We compare three approaches: domain-specific MCP tools (parameterized Cypher templates), text-to-Cypher via the schema-aware NLQ endpoint (GPT-4o with full schema system prompt and few-shot examples), and standalone GPT-4o (no database access).
+The BiomedQA benchmark is open-source.
+[^2] https://github.com/samyama-ai/biomedqa
 Table 10: BiomedQA results (40 questions, 7.9M nodes, 3 federated KGs).
 Approach Accuracy Avg Latency Avg Tokens
-GPT-4o standalone 30/40 (75%) 2,474 ms 213
-Text-to-Cypher (GPT-4o) 0/40 (0%) 986 ms 548
-MCP tools 39/40 (98%) 651 ms 0
-The text-to-Cypher result (0%) is striking: GPT-4o generates syntactically plausible Cypher that silently returns empty results due to schema mismatches (hallucinated property names, incorrect multi-MATCH patterns).
-MCP tools eliminate this class of failure entirely.
-The benchmark and evaluation scripts are open-source.
+GPT-4o standalone 30/40 (75%) 2,805 ms 195
+Text-to-Cypher (NLQ) 34/40 (85%) 1,846 ms 0
+MCP tools 39/40 (98%) 920 ms 0
+Text-to-Cypher achieves 85% with a schema-aware NLQ pipeline (full schema in system prompt, few-shot examples).
+Its 6 failures are: 3 schema hallucinations (non-existent edge traversals), 1 exact-vs-CONTAINS mismatch, 1 inline property variable, and 1 correct empty result.
+MCP tools eliminate schema errors entirely.
+(Token count is 0 because the NLQ endpoint handles the LLM call server-side.)
 7.4 Federation Correctness
-We validated cross-KG federation on the Mac Mini M4 with the following results:
+We validated cross-KG federation on the AWS g4dn.4xlarge with all three KGs loaded (7.9M nodes):
   1. 1.
-    Bridge construction: 32 breast cancer drug nodes were created from a curated DrugBank mapping, yielding 36 TARGETS edges (Drug $\rightarrow$ Protein) and 199 CODED_AS_DRUG edges (Intervention $\rightarrow$ Drug).
-The 1,544 matched drug interventions span 5,338 distinct drug names in breast cancer trials.
+    Drug Interactions $\to$ Pathways bridge: The query “Metformin gene targets $\to$ biological pathways” uses WHERE p.name = g.gene_name to bridge Gene nodes from DGIdb to Protein nodes from Reactome/STRING.
+Returns pathway memberships (e.g., HNF1B $\to$ Developmental Biology) in 3.0 s.
   2. 2.
-    Join overlap: 37,990 Protein nodes with uniprot_id from the Pathways KG are available for joining.
-Of the 36 drug-target proteins, all are present in the Pathways KG’s protein-protein interaction network.
+    Drug Interactions $\to$ Clinical Trials bridge: The query “clinical trials testing Warfarin” uses WHERE i.name = d.name to bridge Drug nodes to Intervention nodes.
+Returns trial NCT IDs and phases (e.g., NCT00835861, PHASE2) in 0.6 s.
   3. 3.
-    Federation query result: The 6-hop query “pathways disrupted by breast cancer trial drugs” returns 10 pathways, with Disease (21 drugs), Signal Transduction (19 drugs), Immune System (12 drugs), and Cell Cycle (12 drugs) as the top results.
-These align with known breast cancer biology: HER2 signaling (Trastuzumab), CDK4/6 cell cycle control (Palbociclib, Ribociclib), and immune checkpoint blockade (Pembrolizumab, Atezolizumab).
+    Three-KG chain: The query “drugs indicated for diabetes $\to$ gene targets $\to$ pathways” traverses Drug $\to$ Indication (Drug Int.
+KG) $\to$ Gene (Drug Int.
+KG) $\to$ Protein $\to$ Pathway (Pathways KG), returning pathways such as Circadian clock and Dissolution of Fibrin Clot via SERPINE1, in 3.9 s.
 8 Interactive Visualization
@@
   3. 3.
-    Memory requirements: The combined 7.8M-node graph requires approximately 8GB RAM. Machines with less memory can load the Pathways KG alone (119K nodes, <<500MB).
+    Memory requirements: The combined 7.9M-node graph requires approximately 33 GB RAM. Machines with less memory can load the Pathways + Drug Interactions KGs alone (151K nodes, <<1 GB).
   4. 4.
-    No cross-tenant queries: Currently, federation requires loading both snapshots into a single tenant. Native cross-tenant query support is planned for a future Samyama release.
+    No cross-tenant queries: Currently, federation requires loading all snapshots into a single tenant. Native cross-tenant query support is planned for a future Samyama release.
 9.3 Generalizability
@@
 Loading all three snapshots into a single graph tenant enables cross-KG federated queries that bridge molecular biology, translational medicine, and pharmacogenomics.
-Cross-KG queries complete in 80 ms–4 s on commodity hardware (62 GB RAM).
+Single-KG queries complete in 80–100 ms; cross-KG federation joins in 1–4 s on commodity cloud hardware (62 GB RAM).
 A Rust native loader constructs the Drug Interactions KG in under 1 second.
 We introduced domain-specific MCP tools for LLM agent access and evaluated them on a new BiomedQA benchmark (40 pharmacology questions): MCP tools achieve 98% accuracy vs.
-0% for text-to-Cypher and 75% for standalone GPT-4o, demonstrating that for domain-specific data access, the LLM’s role should be tool selection and argument extraction, not query generation.
+85% for schema-aware text-to-Cypher and 75% for standalone GPT-4o—with zero schema errors, demonstrating that for domain-specific data access, the LLM’s role should be tool selection and argument extraction, not query generation.
 All code, snapshots, ETL pipelines, benchmarks, and MCP configurations are open-source. Researchers can reproduce the full 3-KG federation—from snapshot import to cross-KG queries—in under 3 minutes.